OpenStudy (anonymous):

solve for r in the equation C=2d-2r

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

I kept getting an answer that wasn't there

OpenStudy (anonymous):

i kept getting (c-2d)/-2=(-2r)/-2 which i simplified to (c/-2)d=r but that cant be right since that wasnt an answer choice

jimthompson5910 (jim_thompson5910):

here is a similar example if you're completely stuck \[\large 2x-3y = 5\] solve for y \[\large 2x-3y = 5\] \[\large -3y = 5-2x\] \[\large y = \frac{5-2x}{-3}\] \[\large y = \frac{-5+2x}{3}\]

OpenStudy (anonymous):

its not that i don't know how to do these types of questions, im actually fairly comfortable with it, i just cant seem to get a "correct" answer; either im right and the question is wrong, or (more likely) the question is legitimate and im wrong and keep making some simple mistake @marissalovescats @phi

OpenStudy (marissalovescats):

Hey. The best of the best make simple mistakes. The more practice you get, the better you will get :) Even I forget stuff like a negative times a negative is a positive. Haha.

OpenStudy (anonymous):

is that what i did wrong? id i forget a sign change somewhere?

OpenStudy (marissalovescats):

No.. I was just showing an example of mistakes I make :P

OpenStudy (marissalovescats):

Whats the 1st step you should do with C=2d-2r to get r alone?

OpenStudy (phi):

i kept getting (c-2d)/-2=(-2r)/-2 which i simplified to (c/-2)d=r the first part is ok \[ \frac{c-2d}{-2} = r\]

OpenStudy (anonymous):

@marissalovescats subtract 2d from both sides which gives me C-2d=-2r

OpenStudy (marissalovescats):

Yup

OpenStudy (phi):

the second part simplified to (c/-2)d=r looks dubious you could say c/-2 + -2d/-2 = c/-2 + d but if we start with \[ \frac{c-2d}{-2} = r \] and don't see this choice. try multiplying -1 top and bottom \[ \frac{-(c-2d)}{2} = r \\ \frac{-c+2d}{2} = r\]

OpenStudy (anonymous):

oh. i feel like i could have totally done that. giant help, everyone!